2(y+z)dx-(x+z)dy+(2y-x+z)dz=0


1
Expert's answer
2021-03-08T15:55:18-0500

Solution:

Given: "2(y+z)dz-(x+z)dy+(2y-x+z)dz=0"

We know that the general form of the Pfaffian equation is

"P d x+Q d y+R d z=0"

The integrability condition for the Pfaffian equation is

"(\operatorname{curl\ F}, F)=0"

where "F=(P, Q, R)" , or

"P\left(\frac{\partial Q}{\partial z}-\frac{\partial R}{\partial y}\right)+Q\left(\frac{\partial R}{\partial x}-\frac{\partial P}{\partial z}\right)+R\left(\frac{\partial P}{\partial y}-\frac{\partial Q}{\partial x}\right)=0"

Here, "P=2(y+z), Q=-(x+z), R=(2 y-x+z)"

Now, verification: LHS

"=2(y+z)(-1-2)-(x+z)(-1-2)+(2 y-x+z)(2+1)"

"=-6 y-6 z+3 x+3 z+6 y-3 x+3 z\\=0"

"=" RHS

Hence, verified.

So, the integrability condition for this equation hold.

If the Pfaffian equation is multiplied by a certain function "\mu(x, y, z)" then one can obtain in the

left-hand side the total differential.

Next, if treating "z" as a constant, then the given equation reduces to

"2(y+z) d x-(x+z) d y=0 \\"

"\Rightarrow\frac{2}{x+z} d x-\frac{1}{y+z} d y=0"

which has a solution

"U(x, y, z)=\frac{(x+z)^{2}}{y+z}=C_{1}"

The integrating factor "\mu(x, y, z)" is given by

"\mu=\frac{1}{P} \cdot \frac{\partial U}{\partial x}=\frac{1}{2(y+z)} \cdot \frac{2(x+z)}{y+z}=\frac{x+z}{(y+z)^{2}}"

Then, we get

"\begin{array}{c}
K=\mu R-\frac{\partial U}{\partial z}= \\
=\frac{x+z}{(y+z)^{2}}(2 y-x+z)-\frac{2(x+z)(y+z)-(x+z)^{2}}{(y+z)^{2}}= \\
=\frac{x+z}{(y+z)^{2}}(2 y-x+z-2 y-2 z+x+z)=0
\end{array}"


Thus we have the equation

"d U=0"

which has a solution "U=c"

Thus the solution of the given equation is

"(x+z)^{2}=c(y+z)"

where "c" is a constant.